# A LoD of Gaussians: Unified Training and Rendering for Ultra-Large Scale Reconstruction with External Memory

### v3

@@ Abstract
1 Introduction
@@
Given a set of posed images of a 3D scene, the task of novel view synthesis (NVS) is to generate plausible images of the scene from unseen viewpoints. Early approaches achieved this via image-based blending (DeepBlending), but the introduction of Neural Radiance Fields (NeRF) (NeRF) marked a breakthrough, enabling high-quality results by optimizing an implicit volumetric scene representation through a multi-layer perceptron. More recently, 3D Gaussian Splatting (3DGS) (3DGS) extended this paradigm to an explicit representation: a set of Gaussian primitives that are efficiently rasterized using splatting techniques (zwicker2001EWA), replacing costly ray marching and allowing real-time rendering with fast convergence.
@@
Despite these advances, both NeRF and 3DGS remain constrained by memory bottlenecks when applied to large-scale environments. Prior methods address this by dividing scenes into smaller chunks (GridNeRF; BlockNeRF; Hierarchical; CityGaussian; GigaGS; VastGaussian), training each independently before merging results. While chunking strategies mitigate memory usage during training, they introduce several key limitations:
  1. 1.
@@
    View-chunk misalignment: Camera views often span multiple chunks, especially in open environments or multi-scale datasets (e.g., combining aerial and street-level images). This makes chunk boundaries arbitrary and hard to define, complicating scene partitioning and training. We identify two artifacts caused by view-chunk misalignment: Chunk bleeding occurs when Gaussians extend out of their current chunk and obscure neighbouring chunks after merging. When an occluder on a training image is not part of the current chunk, its ”ghostly” outline will be trained into the wrong chunk, leading to Chunk ghosting. Examples of these artifacts can be seen in Figures6 and 11.
  2. 2.
@@
    Redundant overlap: To avoid artifacts at chunk boundaries, regions are typically trained with significant overlap, which increases memory usage and prolongs training time.
  3. 3.
@@
    Asymmetric hardware demands: Although chunking reduces memory requirements during training, rendering may require all visible chunks in memory simultaneously—often exceeding the capacity of the original training setup.
@@
The simplest and most robust alternative to chunking is to avoid splitting altogether. With A LoD of Gaussians, we introduce a seamless pipeline that enables training and rendering of ultra-large-scale scenes directly on a single consumer-grade GPU, without any form of scene partitioning (see Figure 7). To handle scenes that exceed available VRAM, we store all Gaussian data in CPU RAM and dynamically stream only those visible from the current training view into GPU memory. Still, a single distant view could require access to the full scene. To address this, we construct a hierarchical Level-of-Detail (LoD) model inspired by Hierarchical, loading detail proportional to view distance. Maintaining this hierarchy during training is non-trivial, as Gaussian properties evolve dynamically. We propose a novel hierarchy densification strategy, adapted from the MCMC-style spawning (3DGSasMCMC), to support stable, progressive refinement. Efficient view-dependent selection from the hierarchy is challenging for large models. Instead of full tree traversal, we adopt Sequential Point Trees (SPTs) (SPT), originally developed for point cloud LoD rendering. Our Hierarchical SPT version allows us to compute the correct LoD cut efficiently for rendering individual views and camera paths. Finally, to reduce CPU-GPU data transfer overhead, we introduce a lightweight caching system that tracks recently used Gaussians and reuses them across training iterations. In summary, we make the following contributions:
  1. 1.
@@
    We propose a novel hierarchy densification strategy that enables dynamic expansion and restructuring of Gaussian LoD representations during training.
  2. 2.
    We leverage the Sequential Point Tree (SPT) data structure and adapt it for large-scale Gaussian Splatting to perform fast, parallelizable LoD cuts for efficient training and rendering.
  3. 3.
    We demonstrate, for the first time, how external memory can be used to train and render ultra-large Gaussian scenes seamlessly on consumer-grade GPUs.
  4. 4.
    We design a caching and view scheduling system that exploits temporal coherence to minimize data transfer overhead and improve training throughput.
  5. 5.
    We verify trough extensive evaluations that the pipeline succeeds in reconstructing huge datasets with complex camera distribution, where SOTA divide-and-conquer methods fail even with larger VRAM budgets.
2 Related Work
@@
Reconstructing large-scale scenes from images has long been a central challenge in visual computing. Traditional approaches relied on Structure-from-Motion (SfM) pipelines to recover geometry from unordered photo collections (RomeInADay; COLMAP). Differentiable rendering techniques, notably NeRF (NeRF) and 3DGS (3DGS), marked a paradigm shift by optimizing volumetric scene representations. Extensions of NeRF to large scenes typically employ scene partitioning (BlockNeRF; GridNeRF) or multi-GPU training strategies (NerfXL). Similarly, most large-scale 3DGS pipelines adopt chunk-based training: Hierarchical-3DGS (H-3DGS) (Hierarchical) trains chunks independently and then merges them into a global LoD hierarchy; CityGaussian (CityGaussian) combines chunked training with per-chunk LoD selection using LightGaussian (LightGaussian); and VastGaussian (VastGaussian) introduces decoupled appearance modeling and progressive partitioning. Horizon-GS (HorizonGS) integrates divide-and-conquer strategies with LoD mechanisms from OctreeGS, specifically targeting hybrid aerial/street-view datasets. GrendelGS (Grendel) avoids spatial chunking by distributing training images across GPUs, such that each device renders a disjoint screen region. Another research direction focuses on extracting geometric proxies from large-scale 3DGS scenes (ULSRGS; CityGaussianV2; GigaGS). These methods leverage TSDF fusion and geometric losses to generate multiple meshes, which are fused and rendered efficiently using traditional rasterization.
@@
Level-of-detail techniques reduce the geometric complexity of distant scene content to accelerate rendering. In the context of 3DGS, LoD approaches have been explored for enabling efficient rendering on memory-constrained or mobile devices. Compression-based strategies include attribute quantization via codebooks, pruning of low-impact Gaussians, and adapting the degree of spherical harmonics per primitive (ReducingTheMemoryFootprint; Compressed3DGaussianSplatting; LightGaussian; MiniSplatting; HierarchicalCompression; RadSplat; FLoD). Scaffold-GS (ScaffoldGS) uses latent vectors anchored to reference Gaussians, with an MLP generating associated Gaussians at render time. Octree-GS (OctreeGS) extends this concept to hierarchical LoD rendering, enabling real-time control over detail levels through spatial subdivision.
3 Preliminaries
Here, we briefly review 3DGS and how Gaussian hierarchies enable adaptive level-of-detail rendering.
3.1 Hierarchical 3D Gaussian Splatting
3DGS (3DGS) models a radiance field using a set of spatially distributed Gaussians, each with mean $\bm{\mu}_{i}\in\mathbb{R}^{3}$, RGB base colors $\mathbf{b}_{i}\in\mathbb{R}^{3}$ and covariance matrices $\bm{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{\top}\mathbf{R}_{i}^{\top}$, which are parameterized via a diagonal scaling matrix $\mathbf{S}_{i}=\text{diag}(s^{1}_{i},s^{2}_{i},s^{3}_{i})$ and an orthonormal rotation matrix $\mathbf{R}_{i}$. Each Gaussian also stores an opacity $\sigma_{i}$ and view-dependent color, modeled using spherical harmonics (SH) coefficients $\mathbf{f}_{i}^{d}$. The SH degree $d$ controls expressiveness, with each Gaussian requiring $\sum_{j=1}^{d}3\cdot(2j+1)$ parameters. For rendering, all $N$ Gaussians are sorted by distance to the camera and a discrete approximation of the volume rendering equation is evaluated for every pixel $\mathbf{x}$ with corresponding view direction $\mathbf{v}$:
$\mathbf{C}(\mathbf{x})=\sum_{i=1}^{N}\mathbf{c}_{i}(\mathbf{v})\alpha_{i}(\mathbf{x})\prod_{j=1}^{i-1}(1-\alpha_{j}(\mathbf{x})),$ (1)
@@ 3.1 Hierarchical 3D Gaussian Splatting
$\alpha_{j}(\mathbf{x})=\sigma_{j}e^{-\frac{1}{2}(\mathbf{x}-\bm{\mu^{\prime}}_{j})\bm{\Sigma^{\prime}}(\mathbf{x}-\bm{\mu^{\prime}}_{j})^{T}}.$ (2)
Here $\bm{\mu^{\prime}}$ and $\bm{\Sigma^{\prime}}$ denote the projected 2D mean and covariance on the image plane, obtained by applying an affine approximation of the projective transform (zwicker2001EWA).
@@
Standard 3DGS pipelines store the full set of per-Gaussian attributes, training images, and optimizer state in GPU memory (VRAM); see Figure A.1 for a detailed per-Gaussian breakdown. Additional temporary allocations occur during forward and backward passes (e.g., for sorting and gradient accumulation). This overhead varies with the scale of the Gaussians and effectiveness of culling strategies, but can be roughly upper-bounded by around 800 bytes per Gaussian in practice. This limits typical training to roughly 500 000500\,000 Gaussians per GB of GPU memory, imposing strict constraints on the detail and extent of reconstructions.
@@
as introduced in H-3DGS (Hierarchical), recursively merge nearby Gaussians into a tree, where each non-leaf node approximates its children, and leaves correspond to the original Gaussians. A cut is defined by a condition $c_{\text{hier}}(i,\text{cam})$ evaluated in a breadth-first search (BFS). If a node satisfies the cut condition, it is added to the cut set and its children are skipped; otherwise, the BFS continues. A proper cut set includes no parents or children of any included node and thus provides a view-adaptive LoD representation.
The cut condition used in Hierarchical is a simple cutoff to the camera distance:
@@
$c_{\text{hier}}(i,\text{cam})=\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i),\quad m_{d}(i)=\frac{T}{\max_{j}s_{i}^{j}},$ (3)
@@
where $\mathbf{p}_{cam}$ is the camera position, $T$ is a global LoD threshold, and $m_{d}(i)$ is the minimum acceptable distance for viewing Gaussian $i$. The BFS ensures that if $i$ is in the cut set, parent($i$) failed the condition: $m_{d}(\text{parent}(i))>\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
4 Method
@@
To train models that exceed GPU memory limits, all Gaussian attributes are stored in CPU RAM and streamed to the GPU on demand for each training view according to the LoD hierarchy. To accelerate the hierarchy cut, we store a copy of only the tree structure in VRAM, where larger subtrees are replaced by Sequential Point Trees (SPTs), forming a hierarchical SPT (HSPT). To minimize costly transfers between RAM and VRAM, we track which SPTs currently reside in GPU memory and at which detail. Only if an SPT is not present in this GPU cache at a similar level of detail, will it be loaded from RAM. Densification is performed on the CPU by adding new leaf nodes to the hierarchy and respawning low-opacity leaf nodes. The densified hierarchy is then converted back to an HSPT and transferred to the GPU for a new round of training iterations. Appendix A.2 includes more details on initialization and training. An overview of our training and densification process can be found in Figure 8 and pseudocodes in Appendix A.8.
4.1 Sequential Point Trees for Gaussian Splatting

### v4

# A LoD of Gaussians: Out-of-Core Training and Rendering for Seamless Ultra-Large Scene Reconstruction

@@ Abstract.
Level of Detail, Gaussian Splatting, Large-Scale Reconstruction
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06submissionid: 395ccs: Computing methodologies Rasterizationccs: Computing methodologies Visibilityccs: Computing methodologies Machine learning
Figure 1. Teaser: We introduce a fully hierarchical 3D Gaussian representation trained directly across unstructured, multi-scale image sets—including street-level and far aerial views—without scene partitioning. Our method maintains a consistent global scene model, eliminating boundary artifacts typical of chunked approaches. A hybrid Level-of-Detail system combines Gaussian hierarchies with Sequential Point Trees, enabling dynamic, view-dependent streaming and LoD selection. The entire model resides in external memory, with only a small, adaptive subset loaded on demand, allowing seamless training and interactive rendering of scenes with 150M+ Gaussians on a single consumer GPU ($\leq 24GB$ VRAM).
1. Introduction
@@
Given a set of posed images of a 3D scene, the task of novel view synthesis (NVS) is to generate plausible images of the scene from unseen viewpoints. Early approaches achieved this via image-based blending (Zhang et al., 2020), but the introduction of Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) marked a breakthrough, enabling high-quality results by optimizing an implicit volumetric scene representation through a multi-layer perceptron. More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) extended this paradigm to an explicit representation: a set of Gaussian primitives that are efficiently rasterized using splatting techniques (Zwicker et al., 2001), replacing costly ray marching and allowing real-time rendering with fast convergence.
@@
Despite these advances, both NeRF and 3DGS remain constrained by memory bottlenecks when applied to large-scale environments. Prior methods address this by dividing scenes into smaller chunks (Xu et al., 2023; Tancik et al., 2022; Kerbl et al., 2024; Liu et al., 2024a; Chen et al., 2025; Lin et al., 2024), training each independently before merging results. While chunking strategies mitigate memory usage during training, they introduce several key limitations:
  1. (1)
@@
    View-chunk misalignment: Camera views often span multiple chunks, especially in open or multi-scale datasets (e.g., combining aerial and street-level images). As a result, chunk boundaries become arbitrary with respect to images, complicating scene partitioning and training, resulting in artifacts: Chunk bleeding occurs when Gaussians extend out of their assigned chunk and obscure neighbouring chunks after merging. Chunk ghosting occurs when an occluder present in a training image is not part of the current chunk, training its ‘ghostly’ outline into the wrong chunk. See Figures 7 and 12.
  2. (2)
@@
    Redundant overlap: To avoid artifacts at chunk boundaries, regions are typically trained with significant overlap, which duplicates parameters and optimizer state, increasing memory usage and prolonging training time.
  3. (3)
@@
    Asymmetric hardware demands: Although chunking reduces memory requirements during training, rendering may require all visible chunks in memory simultaneously—often exceeding the capacity of the original training setup and undermining the practical benefit of partitioning.
@@
The simplest and most robust alternative to chunking is to avoid splitting altogether. With A LoD of Gaussians, we introduce a seamless pipeline that enables training and rendering of ultra-large-scale scenes directly on a single consumer-grade GPU, without any form of scene partitioning (see Figure 7). To handle scenes that exceed available VRAM, we store all Gaussian data in CPU RAM and dynamically stream only those visible from the current training view into GPU memory. Still, a single distant view could require access to the full scene. To address this, we construct a hierarchical Level-of-Detail (LoD) model inspired by Kerbl et al. (2024), loading detail proportional to view distance. Crucially, this hierarchy must be maintained throughout training, as Gaussian parameters and their spatial distribution evolve dynamically. We propose a novel hierarchy densification strategy, inspired by MCMC-style spawning (Kheradmand et al., 2024), to support stable and progressive refinement. Efficient view-dependent selection from the hierarchy is challenging for large models, particularly when the hierarchy structure changes during optimization. Instead of full tree traversal, we adopt Sequential Point Trees (SPTs) (Dachsbacher et al., 2003), originally developed for point cloud rendering. Our Hierarchical SPT version allows us to compute the correct LoD cut efficiently for individual views and camera paths. Finally, to reduce CPU-GPU data transfer, we introduce a lightweight caching system that tracks recently used Gaussians and reuses them across training iterations. In summary, we make the following contributions:
  1. (1)
    Seamless, non-partitioned training of ultra-large Gaussian scenes. We present the first 3D Gaussian Splatting framework that enables training and interactive rendering of city-scale scenes from arbitrary views without spatial partitioning, using out-of-core memory and view-dependent streaming on a single consumer GPU.
  2. (2)
@@
    Dynamic LoD hierarchy densification during training. We propose a novel coarse-to-fine hierarchy densification strategy that allows Gaussian LoD hierarchies to evolve continuously during optimization, supporting stable refinement and restructuring without post-training hierarchy construction.
  3. (3)
    Hierarchical Sequential Point Trees for Gaussian Splatting. We adapt Sequential Point Trees (SPTs) to Gaussian splatting and introduce the Hierarchical SPT (HSPT), a hybrid data structure enabling efficient, parallelizable LoD selection while remaining robust to hierarchy updates during training.
  4. (4)
    Efficient out-of-core execution via cache-aware streaming and view scheduling. We design a lightweight GPU caching and view selection strategy that exploits temporal coherence to substantially reduce CPU–GPU transfer overhead during both training and rendering.
  5. (5)
    Large-scale evaluation and new benchmarks. We introduce the Uni10k dataset and an expanded version of the MatrixCity Small City scene, and demonstrate state-of-the-art performance on large-scale, multi-view datasets spanning aerial, street-level, and indoor environments.
2. Related Work
@@
Reconstructing large-scale scenes from images has long been a central challenge in visual computing. Traditional approaches relied on Structure-from-Motion (SfM) pipelines to recover geometry from unordered photo collections (Agarwal et al., 2009; Schönberger and Frahm, 2016). Differentiable rendering techniques, notably NeRF (Mildenhall et al., 2021) and 3DGS (Kerbl et al., 2023), marked a paradigm shift by optimizing volumetric scene representations. Extensions of NeRF to large scenes typically employ scene partitioning (Tancik et al., 2022; Xu et al., 2023) or multi-GPU training strategies (Li et al., 2024b). Similarly, most large-scale 3DGS pipelines adopt chunk-based training: Hierarchical-3DGS (H-3DGS) (Kerbl et al., 2024) trains chunks independently and then merges them into a global LoD hierarchy; CityGaussian (Liu et al., 2024a) combines chunked training with per-chunk LoD selection using LightGaussian (Fan et al., 2024); OccluGaussian (Liu et al., 2025) partitions the scene to maximize camera correlation in each chunk; and VastGaussian (Lin et al., 2024) introduces decoupled appearance modeling and progressive partitioning. Horizon-GS (Jiang et al., 2025) integrates divide-and-conquer strategies with LoD mechanisms from Ren et al. (2024), specifically targeting hybrid aerial/street-view datasets. GrendelGS (Zhao et al., 2024) avoids spatial chunking by distributing training images across GPUs, such that each device renders a disjoint screen region. Another research direction focuses on extracting geometric proxies from large-scale 3DGS scenes (Li et al., 2025; Liu et al., 2024b; Chen et al., 2025). These methods leverage TSDF fusion and geometric losses to generate multiple meshes, which are fused and rendered efficiently using traditional rasterization. Other approaches (Zhao et al., 2026; Lee et al., 2025) make use out-of-core memory to scale training, but neglect LoD, which limits their memory reduction during training to frustum culling.
@@
Level-of-detail techniques reduce geometric complexity of distant scene content to accelerate rendering. In 3DGS, LoD approaches have mainly targeted efficient rendering on memory-constrained or mobile devices. Compression-based strategies include attribute quantization via codebooks, pruning low-impact Gaussians, and adapting the degree of spherical harmonics per primitive (Papantonakis et al., 2024; Niedermayr et al., 2024; Fan et al., 2024; Fang and Wang, 2024; Huang et al., 2025; Niemeyer et al., 2025; Seo et al., 2024). Scaffold-GS (Lu et al., 2024) uses latent vectors anchored to reference Gaussians, with an MLP generating associated Gaussians at render time. Octree-GS (Ren et al., 2024) extends this to hierarchical LoD rendering via spatial subdivision. Virtualized 3D Gaussians (Yang et al., 2025) targets composed scenes of reconstructed objects using an LoD scheme inspired by Unreal Engine 5’s Nanite (Karis et al., 2021).
3. Preliminaries of Hierarchical 3D Gaussian Splatting
3DGS (Kerbl et al., 2023) models a radiance field using a set of spatially distributed Gaussians, each with mean $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, RGB base colors $\mathbf{b}_{i}\in\mathbb{R}^{3}$ and covariance matrices $\boldsymbol{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{\top}\mathbf{R}_{i}^{\top}$, which are parameterized via a diagonal scaling matrix $\mathbf{S}_{i}=\text{diag}(s^{1}_{i},s^{2}_{i},s^{3}_{i})$ and an orthonormal rotation matrix $\mathbf{R}_{i}$. Each Gaussian also stores an opacity $\sigma_{i}$ and view-dependent color, modeled using spherical harmonics (SH) coefficients $\mathbf{f}_{i}^{d}$. The SH degree $d$ controls expressiveness, with each Gaussian requiring $\sum_{j=1}^{d}3\cdot(2j+1)$ parameters. For rendering, all $N$ Gaussians are sorted by distance to the camera and a discrete approximation of the volume rendering equation is evaluated for every pixel $\mathbf{x}$ with corresponding view direction $\mathbf{v}$:
(1) $\mathbf{C}(\mathbf{x})=\sum_{i=1}^{N}\mathbf{c}_{i}(\mathbf{v})\alpha_{i}(\mathbf{x})\prod_{j=1}^{i-1}(1-\alpha_{j}(\mathbf{x})),$
@@ 3. Preliminaries of Hierarchical 3D Gaussian Splatting
(2) $\alpha_{j}(\mathbf{x})=\sigma_{j}e^{-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu^{\prime}}_{j})\boldsymbol{\Sigma^{\prime}}(\mathbf{x}-\boldsymbol{\mu^{\prime}}_{j})^{T}}.$
Here $\boldsymbol{\mu^{\prime}}$ and $\boldsymbol{\Sigma^{\prime}}$ denote the projected 2D mean and covariance on the image plane, obtained by applying an affine approximation of the projective transform (Zwicker et al., 2001).
@@
Standard 3DGS pipelines store the full set of per-Gaussian attributes, training images, and optimizer state in GPU memory (VRAM); see Figure 9 for a detailed per-Gaussian breakdown. Additional temporary allocations occur during forward and backward passes (e.g., for sorting and gradient accumulation). This overhead varies with the scale of the Gaussians and effectiveness of culling strategies, but can be roughly upper-bounded by around 800 bytes per Gaussian in practice. This limits typical training to roughly 500 000500\,000 Gaussians per GB of GPU memory, imposing strict constraints on the detail and extent of reconstructions.
@@
As introduced in H-3DGS (Kerbl et al., 2024), Gaussian hierarchies recursively merge nearby Gaussians into a tree, where each non-leaf node approximates its children, and leaves correspond to the original Gaussians. A cut is defined by a condition $c_{\text{hier}}(i,\text{cam})$ evaluated in a breadth-first search (BFS). If a node satisfies this condition, it is added to the cut set and its children are skipped; otherwise, the BFS continues. A proper cut set includes no parents or children of any included node and thus provides a view-adaptive LoD representation.
The cut condition used in (Kerbl et al., 2024) is a simple cut-off according camera distance:
@@
(3) $c_{\text{hier}}(i,\text{cam})=\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i),\quad m_{d}(i)=\frac{T}{\max_{j}s_{i}^{j}},$
@@
where $\mathbf{p}_{cam}$ is the camera position, $T$ is a global LoD threshold, and $m_{d}(i)$ is the minimum acceptable distance for viewing Gaussian $i$. The BFS ensures that if $i$ is in the cut set, parent($i$) failed the condition: $m_{d}(\text{parent}(i))>\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
4. Method
@@
To train models that exceed GPU memory limits, all Gaussian attributes are stored in CPU RAM and streamed to the GPU on demand for each training view according to the LoD hierarchy. To accelerate the hierarchy cut, we store a copy of only the tree structure in VRAM, where larger subtrees are replaced by Sequential Point Trees (SPTs), forming a hierarchical SPT (HSPT). To minimize costly transfers between RAM and VRAM, we track which SPTs currently reside in GPU memory and at which detail. Only if an SPT is not present in this GPU cache at a sufficiently similar level of detail, will it be loaded from RAM. Densification is performed on the CPU by adding new leaf nodes to the hierarchy and respawning low-opacity leaf nodes. Following densification, the updated hierarchy is converted back to an HSPT and transferred to the GPU for a new round of training iterations. Suppl. A.2 includes more details on initialization and training. An overview of our training and densification process can be found in Figure 9 and pseudocode in Suppl. A.7.
4.1. Sequential Point Trees for Gaussian Splatting
(a) Hierarchy cuts
(b) Densification example